OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (1, 9); Slope: 5 y – 1 = 5(x + 9) y – 9 = 5(x + 1) y – 9 = 5(x – 1) y + 9 = 5(x – 1)

OpenStudy (anonymous):

use this equation y-y1 = m(x-x1) here x1 = 1 y1 = 9 and m = 5

OpenStudy (anonymous):

ok so what i had at first is correct the answer would be C